# RDF Stream Taxonomy: Systematizing RDF Stream Types in Research and Practice

### v2

@@ 4 RDF-STaX Ontology
@prefix rdfs: <http://www.w3.org/2000/01/rdf-schema#> .
@@
@prefix stax: <https://w3id.org/stax/ontology#> .
@@
#  other properties of the dataset
stax:hasStreamTypeUsage [
a stax:RdfStreamTypeUsage ;
stax:hasStreamType stax:datasetStream ;
rdfs:comment "The data is a sequence of RDF datasets."@en
%**** main.tex Line 350 ****a stax:RdfStreamTypeUsage ;
stax:hasStreamType stax:flatQuadStream ;
rdfs:comment "The data can be viewed as a flat sequence of RDF quads."@en
The ontology’s documentationfootnote 3 contains a detailed description of every class, property, and individual.
4.1 Availability and Sustainability
The sources of the ontology and its documentation are hosted on GitHub [^5] https://github.com/RDF-STaX/rdf-stax.github.io. The repository includes continuous integration and deployment (CI/CD) scripts that automatically validate the ontology’s profile, generate documentation, run an OWL 2 reasoner to add more assertions, package the ontology into a number of formats, and publish it on the website [^6] https://w3id.org/stax. Two versions of the ontology are published: OWL 2 DL, which can be used with reasoners; and OWL 2 Full, which contains more metadata and alignments to other vocabularies.
The ontology uses best findability, accessibility, interoperability, and reusability (FAIR) practices developed by the community [22], achieving 86% score in FOOPS! [21] and 4/4 stars in DBpedia Archivo [19]. All releases of the ontology, including the OWL 2 DL and OWL 2 Full variants, are distributed through the RDF-STaX website, under a permanent URL. The ontology is also archived in Zenodo [51] and registered in DBpedia Archivo and Linked Open Vocabularies [61] (pending acceptance).
RDF-STaX is an open and freely-licensed project, welcoming contributions to the ontology, its documentation, and tooling. A usage guidefootnote 4 with several use case scenarios and examples was prepared to facilitate adoption. To ensure long-term sustainability, a comprehensive contribution guide was provided [^7] https://w3id.org/stax/1.0.1/contributing, along with a public issue tracker [^8] https://github.com/RDF-STaX/rdf-stax.github.io/issues. The project is entirely hosted on permanently free infrastructure, which should ensure its long-term stability.
4.2 Comparison with Other Ontologies
To the best of our knowledge, there is no other ontology or vocabulary that would model RDF stream types. As previously noted, the RDF-STaX ontology purposefully does not define classes for RDF streams, but rather focuses on subjective RDF stream usages. This difference makes other ontologies non-overlapping – instead, RDF-STaX can naturally complement them. To this end, the documentation includes examples of using RDF-STaX with these complementary ontologies (VoCaLS, VoID, and DCAT)footnote 4. The usage pattern is to define an RDF stream / dataset with these ontologies as usual, and add the stax:hasStreamTypeUsage property to indicate the stream type, according to RDF-STaX. The same pattern can be used with any future ontology for describing streams or datasets.
The VoCaLS vocabulary for describing published RDF streams [59] defines the RDFStream class for potentially infinite sequences of RDF graphs and/or triples. These streams can be thus annotated as RDF graph streams or flat RDF triple streams in RDF-STaX. The VoID vocabulary [65] describes datasets with RDF data (not RDF datasets as per RDF 1.1), which can be viewed as flat RDF triple streams. The more general DCAT [6] describes datasets and their distributions, which may be RDF streams.
The RDF-STaX ontology also defines four element types (RDF graphs, triples, quads, and datasets). In this regard, we could identify only the SPARQL 1.1 Service Description vocabulary [62], which has the sd:Graph and sd:Dataset classes for RDF graphs and datasets. These classes were aligned with the corresponding terms in RDF-STaX, using the skos:relatedMatch property.
5 Applications
RDF-STaX was created with several applications in mind. Firstly, it can be used to describe stream types used in research and software, to easily compare different solutions and foster compatibility. Secondly, it may be employed to semantically annotate published datasets and streams, improving their accessibility and interoperability. Both of these use cases have already been realized in practice.
5.1 Nanopublications
A nanopublication is a small unit of scientific knowledge published in RDF [31] – it can be an opinion, a measurement, or any other assertion. Nanopublications are a natural way to express statements about scientific works or software due to their strong provenance mechanisms and citeability. This fits well the pattern of subjective statements about stream types, which is used in RDF-STaX. To make it easier to create nanopublications with the RDF-STaX ontology, a template was prepared for Nanodash, an open nanopublication editor [32]. The template along with the accompanying manual [^9] https://w3id.org/stax/1.0.1/nanopubs makes it easy to create and publish semantic assertions such as “X uses stream type Y because Z”, requiring the user only to fill out a simple online form.
The results from the survey conducted in this paper (as summarized in Section 3.3) were published as 32 nanopublications, including a more detailed discussion of the usages than what could fit in the paper. The RDF-STaX nanopublications are packaged as a single knowledge graph in a CI/CD pipeline and published automatically on the websitefootnote 9. Anyone can contribute to the dataset by simply publishing nanopublications with the provided template. The dataset can serve as a living state-of-the-art review, allowing one to easily compare and discuss works on RDF streaming, in a structured manner.
5.2 RiverBench
RiverBench is an open RDF streaming benchmark suite, containing heterogeneous streaming datasets that are published using automated pipelines, following FAIR principles [52]. RiverBench integrates with the RDF-STaX ontology, requiring all datasets to be annotated with their stream types. Each dataset has two stream types assigned, one for the grouped stream formulation and another for the flat. The consistency of the annotations is checked in CI/CD using SHACL rules and the semantic relations defined in RDF-STaX (SKOS taxonomy properties and stax:canBeFlattenedInto). The datasets are then organized into benchmark profiles by their stream type. The published dataset documentation and RDF metadata based on DCAT [6] also includes the RDF stream types.
At the time of writing, RiverBench’s 12 datasets use each stream type in RDF-STaX, except the RDF named graph stream (but there are examples of the timestamped variant). Detailed information about this use case along with examples are available on the website [^10] https://w3id.org/stax/1.0.1/uses.
6 Discussion and Future Work
Although the “RDF stream” is a term that is now ubiquitous in research, it is used to describe many different concepts. There is no universally accepted language for describing RDF streams, which leads to a situation in which RDF streaming solutions become hard to assess, compare, and connect together. For example, forming a pipeline composed of a stream annotator, translator, streaming protocol, and a reasoner currently appears to be a tough challenge, as it is hard to determine what exactly is an RDF stream for each of these tools.
RDF-STaX is embodied in an ontology that makes the taxonomy readily usable in practice. The ontology does not attempt to replace any existing vocabulary for describing streams or datasets, but rather aims to complement them. We hope that by being non-overlapping by design and focused on a single purpose, RDF-STaX will remain useful and relevant in the long term. The two presented use cases, both of which were already realized in practice, are very common and well-known to the community, making the ontology all the more relevant.
Another result of this work is the nanopublications dataset, which is a collaborative, living state-of-the-art review of RDF streaming. Nanopublications, in our view, are a natural way to structure intersubjective scientific discourse. It is possible for example, to comment on other nanopublications, confirm or disagree with them, all in a structured manner. Here we propose to use this powerful tool for a very narrow field of science, but its possible future applications are certainly much broader. We encourage other researchers to contribute their perspectives to the dataset or to comment on the existing assertions.
As for future work, the theory presented in this contribution notably does not tackle the formal semantics of RDF streams. As with RDF datasets, it is unlikely that there can be single, unifying semantics for streams, but nonetheless, we see this area as worth exploring in the future. The presented survey of RDF streaming is also not a systematic review – it only serves as a basis for the discussion in this work. Producing a systematic review of RDF streaming would be very valuable to the community, and we hope that the proposed taxonomy and nanopublication tooling will make creating it easier.
7 Conclusion
In this work we propose the RDF Stream Taxonomy, which systematizes RDF stream definitions found in the literature. The taxonomy uncovered the variety of RDF streaming in research and practice, highlighting the open challenge of making the different applications compatible with each other. The proposed RDF-STaX ontology embodies the semantic relations between stream types in the taxonomy, and makes the definitions readily applicable. The ontology positions itself as complementary to existing vocabularies, and employs a flexible approach of making subjective statements about the types of RDF streams. The ontology and its documentation are fully open and follow best practices for FAIR developed by the community.
The ontology was applied in two use cases – annotating research works in nanopublications, and describing RDF stream types of streaming datasets. The nanopublications dataset is a living state-of-the-art review that anyone can contribute to, using the provided documentation and tooling. The second use case highlights how RDF-STaX can be used to annotate published streaming datasets, integrating with other vocabularies.

### v3

@@ 4.1 Ontology Usage Patterns
[breaklines]turtle @prefix dcat: <http://www.w3.org/ns/dcat#> . @prefix rdfs: <http://www.w3.org/2000/01/rdf-schema#> . @prefix stax: <https://w3id.org/stax/ontology#> .
_:dataset a dcat:Dataset ; # … other properties of the dataset … stax:hasStreamTypeUsage [ a stax:RdfStreamTypeUsage ; stax:hasStreamType stax:datasetStream ; rdfs:comment "The data is a sequence of RDF datasets."@en ] , [ a stax:RdfStreamTypeUsage ; stax:hasStreamType stax:flatQuadStream ; rdfs:comment "The data can be viewed as a flat sequence of RDF quads."@en ] . In some situations it may be more convenient to annotate existing resources from an external perspective. RDF-STaX also supports this, through the use of the stax:isUsageOf property. This is useful, for instance, when annotating published research papers or software with the RDF stream types that they use. This is illustrated in the example below, where the publication about IPSM Ganzha et al. (2017) is annotated:
{minted}
@@
[breaklines]turtle @prefix rdfs: <http://www.w3.org/2000/01/rdf-schema#> . @prefix stax: <https://w3id.org/stax/ontology#> .
_:usage a stax:RdfStreamTypeUsage ; rdfs:comment "The authors state that for their proposed system, ’the streaming element (i.e. a single message) (…) is a set of triples’. Therefore, internally it uses an RDF graph stream."@en ; stax:hasStreamType stax:graphStream ; stax:isUsageOf <https://doi.org/10.1109/ICSTCC.2017.8107003> .
4.2 Alignments to Other Ontologies
The RDF-STaX ontology is aligned to other existing vocabularies with the use of semantic relations to foster the reuse of multiple ontologies linked together. Additionally, alignments can make it easier to understand the concepts introduced in RDF-STaX by comparing them to concepts already present in other vocabularies. However, as detailed in the comparison with the state of the art (Section 6.4), we could identify no other ontologies that would define the same concepts as RDF-STaX in a manner that would enable creating semantically strict, direct alignments (e.g., a subclassing relation). Instead, in cases where other ontologies define terms that are similar to those in RDF-STaX, a skos:relatedMatch alignment is provided. This essentially states that the meaning of the two concepts in different vocabularies is similar.
The following list enumerates the alignments that are provided in RDF-STaX release 1.1.1. Section 6.4 contains further discussion on this topic, comparing the meaning of RDF-STaX’s terms and those of other ontologies.
  • SPARQL 1.1 Service Description vocabulary Williams (2013) – the classes for RDF graphs (sd:Graph) and datasets (sd:Dataset) were aligned with the corresponding terms in RDF-STaX (stax:graph and stax:dataset, respectively), which are instances of class stax:RdfElementType.
  • VoCaLS Tommasini et al. (2018) – the vocals:RDFStream class was aligned with both the stax:flatTripleStream and stax:graphStream instances in RDF-STaX.
  • LDES Van Lancker et al. (2021) – the ldes:EventStream class was aligned with the stax:subjectGraphStream instance in RDF-STaX.
  • VoID Zhao et al. (2011) – the void:Dataset class was aligned with the stax:flatTripleStream instance in RDF-STaX.
4.3 Availability and Sustainability
The sources of the RDF-STaX ontology and its documentation are hosted on GitHub (accessed on 4 June 2024). The repository includes continuous integration and deployment (CI/CD) scripts that automatically validate the ontology’s profile, run competency question tests (see Section 5), generate documentation, run an OWL 2 reasoner to add more assertions, package the ontology into a number of formats, and publish it on the website (accessed on 4 June 2024). Two versions of the ontology are published: OWL 2 DL, which can be used with reasoners (containing 249 triples); and OWL 2 Full (294 triples), which contains more metadata and alignments to other vocabularies. The latter is intended for use in contexts where ontological reasoning is not needed, e.g., for SPARQL querying. The additional metadata present there violates the OWL 2 DL profile, hence producing an OWL 2 Full ontology.
The ontology uses best practices for findability, accessibility, interoperability, and reusability (FAIR) developed by the community Garijo and Poveda-Villalón (2020), compliance with which is evaluated in Section 6.3. All releases of the ontology, including the OWL 2 DL and OWL 2 Full variants, are distributed through the RDF-STaX website, under a permanent URL hosted by the W3C Permanent Identifier Community Group. The ontology is also archived in Zenodo Sowiński (2024) and registered in DBpedia Archivo Frey et al. (2020) (accessed on 4 June 2024) and Linked Open Vocabularies Vandenbussche et al. (2017) (accessed on 4 June 2024).
RDF-STaX is an open and freely-licensed project, welcoming contributions to the ontology, its documentation, and tooling. A usage guide (accessed on 4 June 2024) with several use case scenarios and examples was prepared to facilitate adoption. To ensure long-term sustainability, a comprehensive contribution guide (accessed on 4 June 2024) was provided, along with a public issue tracker (accessed on 4 June 2024). The project is entirely hosted on the permanently free infrastructure of GitHub and w3id.org, which should ensure its long-term stability.
5 Use Cases
In what follows, the three most representative use cases envisioned for RDF-STaX are described. The first use case is annotating published datasets and streams (for example, on the Web), to improve their accessibility and interoperability. The second use case focuses on achieving RDF streaming tool interoperability by embedding stream type metadata within the stream itself. The final envisioned application is analyzing the state of the art (SOTA) of RDF streaming in research and software, to easily compare different solutions and foster compatibility.
All of these use cases are described in detail below, along with accompanying competency questions (CQs) realized as SPARQL queries that should be run against the RDF-STaX ontology in the OWL 2 Full version. The CQ queries are available in RDF-STaX’s documentation (accessed on 4 June 2024) and are used to automatically and continuously test the ontology’s validity with regard to the use cases, as described in Section 6.1. Two of the presented use cases (published dataset annotation and analyzing the SOTA of RDF streaming) are already realized in practice, as detailed below. The three presented use cases should only be considered a starting point for RDF-STaX, as the community may over time find new ways in which to use it. The resource is expected to evolve over time (through a process called ontology evolution in the literature Zablith et al. (2015)) to remain up-to-date with the community’s needs. To enable this, RDF-STaX adopts a thoroughly open paradigm, inviting outside contributions.
5.1 Annotating Published Datasets and Streams
In this use case, RDF-STaX’s stream types are used as metadata of datasets or streams published on the Web (for example, using the Linked Data mechanisms). This use case can be viewed from two sides – the publisher’s, and the end user’s, which is reflected in the defined competency questions (see Table 2). The questions pertain to obtaining basic information about the defined stream types (CQ1.1 and CQ1.2). This can be used by the publisher to decide which stream type to use, and by the end user to understand what that stream type really is. Questions CQ1.3 and CQ1.5 play an analogous role, but also aim to provide additional context by relating the terms in RDF-STaX to external sources (other ontologies). Finally, CQ1.4 serves the end user by providing links to examples of how each stream type can be used in practice.
Table 2: Competency questions for use case 1 – annotating published datasets and streams. Full SPARQL queries and expectations for each test can be found on the ontology’s website (accessed on 4 June 2024).
# Competency question
CQ1.1 What are the names and descriptions of all RDF stream types?
CQ1.2 What is the definition for each stream type?
CQ1.3 What is the type of element of each concrete stream type? Provide additional references to external sources for each element type, if available.
CQ1.4 How can each of the concrete stream types be used? Provide a link to one example for each.
CQ1.5 What are the corresponding terms from other ontologies to those defined in the RDF-STaX ontology?
This use case was implemented in RiverBench Sowiński et al. (2023), an open RDF streaming benchmark suite, containing heterogeneous streaming datasets that are published using automated pipelines, following FAIR principles. RiverBench integrates with the RDF-STaX ontology, requiring all datasets to be annotated with their stream types. Each dataset has two stream types assigned, one for the grouped stream formulation and another for the flat. The consistency of the annotations is checked in CI/CD using SHACL rules Knublauch and Kontokostas (2017) and the semantic relations defined in RDF-STaX (SKOS taxonomy properties and stax:canBeFlattenedInto). The datasets are then organized into benchmark profiles by their stream type. The published dataset documentation and RDF metadata based on DCAT Browning et al. (2023) also includes the RDF stream type information.
At the time of submission, RiverBench’s 2.0.0 release includes 12 datasets that use each stream type in RDF-STaX, except the RDF named graph stream (however, there are examples of the timestamped variant). The full metadata dump of RiverBench 2.0.0 consists of 9280 triples, including 201 stax:hasStreamTypeUsage annotations. Detailed technical information about this use case along with examples are available on the RDF-STaX website (accessed on 4 June 2024).
5.2 Embedding Metadata in Streams
This use case focuses on embedding stream type metadata within the stream itself, with the goal of enabling streaming system interoperability. For example, a stream producer would state that the stream is of a given type, and then the consumer would check if it can consume this type of stream or convert it into a consumable stream. As this use case focuses on automated handling of streams, the competency questions (see Table 3) exploit the reasoning capabilities of RDF-STaX. Thus, CQ2.1, CQ2.2, CQ2.3, and CQ2.4 all address different possible transformations a stream processor could perform on an RDF stream.
Table 3: Competency questions for use case 2 – embedding metadata in streams. Full SPARQL queries and expectations for each test can be found on the ontology’s website (accessed on 4 June 2024).
# Competency question
CQ2.1 Which concrete stream types can be viewed as generalizations of other stream types?
CQ2.2 Which concrete stream types can be trivially extended (by assuming the default graph) into other stream types?
CQ2.3 Which concrete stream types can be flattened into other stream types?
CQ2.4 Which concrete stream types can be grouped to obtain other stream types?
5.3 Analyzing the State of the Art of RDF Streaming
The final use case pertains to analyzing the SOTA of RDF streaming by annotating published research works and software with the RDF stream types used by them. This task is thus very closely related to the SOTA analysis performed in this paper, in Section 3.3. As with any other scientific field, however, the landscape of RDF streaming changes rapidly and therefore using the RDF-STaX ontology in combination with a knowledge base of research works can help automate some literature review processes. Therefore, the defined competency questions (see Table 4) tackle more complex issues that may be asked when analyzing the correspondences between the annotated research works. CQ3.1 and CQ3.2 pertain to the taxonomical structure of RDF-STaX, helping the user understand how the different stream types relate to each other. Meanwhile, CQ3.3 attempts to find pairs of essentially incompatible RDF stream formulations, i.e., those that cannot be trivially converted into each other. This would serve to identify methods that cannot easily interoperate.
Table 4: Competency questions for use case 3 – analyzing the state of the art of RDF streaming. Full SPARQL queries and expectations for each test can be found on the ontology’s website (accessed on 4 June 2024).
# Competency question
CQ3.1 What are the taxonomical parents of each stream type, listed in order?
CQ3.2 Is there any stream type that is a taxonomical parent or child of itself?
CQ3.3 Which conversions between concrete stream types cannot be done in any trivial way (grouping, flattening, extending)? Allow for multiple trivial transformations in series and take into account the taxonomical structure.
This use case was implemented within the scope of this work, with a nanopublication dataset of research works on RDF streaming. A nanopublication is a small unit of scientific knowledge published in RDF Kuhn et al. (2013) – it can be an opinion, a measurement, or any other assertion. Nanopublications are a natural way to express statements about scientific works or software due to their strong provenance mechanisms and citeability. This fits well the pattern of subjective statements about stream types, which is used in RDF-STaX. Therefore, to make it easier to create nanopublications with the RDF-STaX ontology, a template was prepared for Nanodash, an open nanopublication editor Kuhn et al. (2021). The template along with the accompanying manual (accessed on 4 June 2024) makes it easy to create and publish semantic assertions such as “X uses stream type Y because Z”, requiring the user to only fill out a simple online form.
The results from the survey conducted in this paper (as summarized in Section 3.3) were published as 35 nanopublications using the aforementioned mechanism, totaling 876 triples. The nanopublications include a more elaborated discussion of each stream type usage, providing more details than Table 1. The RDF-STaX nanopublications are packaged as a single knowledge graph in a CI/CD pipeline and published automatically on the website (accessed on 4 June 2024), where they can be downloaded and reused under the free CC BY 4.0 license. Anyone can contribute to the dataset by simply publishing nanopublications with the provided Nanodash template. Effectively, the dataset along with the publishing mechanism can serve as a collaborative, living state-of-the-art review, allowing one to easily compare and discuss works on RDF streaming, in a structured manner. A similar mechanism for living literature reviews was previously proposed by Wijkstra et al. Wijkstra et al. (2021), to whose work we refer the reader for more context.
6 Ontology Evaluation
This section presents the results of a multifaceted evaluation of the RDF-STaX ontology. Firstly, the evaluation includes checking use case coverage with the use of competency questions defined in Section 5. Secondly, the validity of the ontology is evaluated with regard to the OWL specification and its logical consistency. Thirdly, the findability, accessibility, interoperability, and reusability aspects of the ontology are evaluated, especially with regard to the current, community-developed best practices. Finally, RDF-STaX is compared with other similar resources, highlighting the advances offered by this work.
6.1 Use Case Coverage
This evaluation is based on the twelve competency question tests defined in Section 5. Each of these tests was written down as a SPARQL query paired with the expectation for the test, which specifies how many results should the query return when run against the RDF-STaX ontology. The tests are run in sequence by a CI/CD job implemented in the Python language, with the rdflib library RDFLib contributors – source code is available on GitHub (accessed on 4 June 2024) under the Apache 2.0 license. The job is triggered on every ontology release (including the development release) and pull request. If any test fails, the ontology will not be released, thus ensuring that it always covers the competency questions from the use cases. Further technical implementation details on the tests are available in the documentation (accessed on 4 June 2024).
At the time of submission (RDF-STaX release 1.1.1), the CI/CD job reports that the ontology successfully passes all competency question tests. Additionally, the fact that the ontology was employed in two implemented applications (see Section 5) serves as additional validation of the ontology’s usefulness and practicality Sowiński et al. (2022).
6.2 Logical and OWL Profile Validity
The purpose of this evaluation is to check if the ontology meets the requirements of specific OWL 2 profiles Fokoue et al. (2012), which strike different balances between the expressive power of the ontology and reasoning efficiency. Similarly to the use case coverage evaluation, the compliance checks with OWL 2 profiles were implemented as CI/CD jobs, which run on any ontology release and pull request. Both distributions of the ontology (OWL 2 DL and OWL 2 Full) are checked to see if they match their respective profile. The check is performed using ROBOT’s Jackson et al. (2019) validate-profile command. Additionally, the logical consistency of the ontology is validated in CI/CD with ROBOT’s reason command, using the HermiT reasoner Glimm et al. (2014).
At the time of submission (RDF-STaX release 1.1.1), the CI/CD job reports that both distributions of the ontology (OWL 2 DL and OWL 2 Full) are compliant with their respective profiles, with no inconsistencies. For additional validation, RDF-STaX 1.1.1 was checked with the publicly available Ontology Pitfall Scanner (OOPS!) Poveda-Villalón et al. (2014), which reported no critical issues.
6.3 Findability, Accessibility, Interoperability, and Reusability
The question of what constitutes a FAIR ontology in practice is a complicated one Poveda-Villalón et al. (2020), with many possible interpretations for each of the FAIR Guiding Principles Wilkinson et al. (2016). To best reflect the needs of the Semantic Web community and ensure impartiality, in this evaluation we chose to rely on two well-known, publicly available, automatic FAIR evaluators: FOOPS! Garijo et al. (2021) and DBpedia Archivo Frey et al. (2020). The following results can be easily reproduced by inputting into either of them the RDF-STaX ontology IRI.
In FOOPS!, RDF-STaX obtained a score of 95%, with 8.83/9 points in findability, 3/3 in accessibility, 3/3 in interoperability, and 7.92/9 in reusability (Figure 3(a)). All of the issues reported by the tool are minor. In DBpedia Archivo, RDF-STaX obtained 4/4 stars (Figure 3(b)). It should be stressed here that Archivo’s manual states that such a result means the ontology has reached “minimum viability” in the context of FAIR but is not necessarily “good”. However, at the time of submission this is the highest possible rating offered by DBpedia Archivo, with only 20.03% of all indexed ontologies having four out of four stars.
(a) FAIR evaluation results obtained from FOOPS! Garijo et al. (2021).
(b) FAIR evaluation results obtained from DBpedia Archivo Frey et al. (2020).
Figure 3: FAIR evaluation results for RDF-STaX. Both screenshots were taken on 4 June 2024.
6.4 Comparison to the State of the Art
To the best of our knowledge, there is no other ontology or vocabulary that would model RDF stream types. As previously noted, the RDF-STaX ontology purposefully does not define classes for RDF streams, but rather focuses on subjective RDF stream usages. This difference makes other ontologies non-overlapping – instead, RDF-STaX can naturally complement them. As shown below, each existing ontology also covers only one or two selected formulations of the RDF streaming problem. Thus, RDF-STaX can provide a common semantic bridge to describe RDF stream types across different underlying vocabularies. To this end, the documentation includes examples (accessed on 4 June 2024) of using RDF-STaX with these complementary ontologies (VoCaLS, VoID, LDES, and DCAT). The expected usage pattern is to define an RDF stream or a dataset with these ontologies as usual, and add the stax:hasStreamTypeUsage property to indicate the stream type according to RDF-STaX. The same pattern can be used with any future ontology for describing streams or datasets.
In what follows, the relevant ontologies are discussed in detail, focusing on their differences and complementarities with RDF-STaX. In the discussion, only publicly available ontologies are included, as otherwise it would not be possible to reliably assess the ontology’s scope and semantics.
The VoCaLS vocabulary Tommasini et al. (2018) is intended to help publish streaming data using the Linked Data principles and to describe streaming services. It defines the vocals:RDFStream class for potentially infinite sequences of RDF graphs and/or triples. Instances of this class can be thus annotated as RDF graph streams or flat RDF triple streams in RDF-STaX. As RDF-STaX does not define a class for RDF streams (but rather has instances of RDF stream types), the two ontologies are entirely complementary. Due to the different semantics of the terms in VoCaLS and RDF-STaX, a direct semantic connection (e.g., subclassing) between the two ontologies cannot be made. Instead, a skos:relatedMatch alignment is provided, along with a suggested usage pattern in the documentation of RDF-STaX.
The Linked Data Event Streams (LDES) vocabulary Van Lancker et al. (2021) defines the ldes:EventStream class, which is a collection of stream elements – sets of RDF triples. Each element is identified with its subject node. Therefore, instances of this class can be annotated in RDF-STaX as RDF subject graph streams. Such an alignment was added to the ontology.
The VoID vocabulary Zhao et al. (2011) describes datasets with RDF data (not RDF datasets as per RDF 1.1), which are in fact sets of triples with additional context information. Thus, instances of class void:Dataset can be annotated as RDF-STaX’s flat RDF triple streams. This alignment was added to the ontology.
The more general DCAT Browning et al. (2023) vocabulary describes datasets and their distributions. These datasets may be RDF streams, but DCAT does not assume any specific format of the datasets annotated with it. Therefore, RDF-STaX can be used as a complementary tool to annotate the type of the stream contained within such a dataset – this is done for example in RiverBench, see Section 5.1.
The RDF-STaX ontology also defines four element types (RDF graphs, triples, quads, and datasets). In this regard, we could identify only the SPARQL 1.1 Service Description vocabulary Williams (2013), which has the sd:Graph and sd:Dataset classes for RDF graphs and datasets. These classes were aligned with the corresponding terms in RDF-STaX, using the skos:relatedMatch property.
In summary, to the best of our knowledge, RDF-STaX is the first taxonomy and ontology that allows describing more than one type of RDF stream, making it a crucial step in enabling RDF streaming interoperability and understanding the field’s state of the art.
7 Discussion and Future Work
Although the “RDF stream” is a term that is now ubiquitous in research, it is used to describe many different concepts. There is no universally accepted language for describing RDF streams, which leads to a situation in which RDF streaming solutions become hard to assess, compare, and connect together. For example, forming a pipeline composed of a stream annotator, translator, streaming protocol, and a reasoner currently appears to be a tough challenge, as it is hard to determine what exactly is an RDF stream for each of these tools and whether these formulations are compatible with each other or not.
RDF-STaX is embodied in an ontology that makes the taxonomy readily usable in practice. The ontology does not attempt to replace any existing vocabulary for describing streams or datasets, but rather aims to complement them and serve as a shared semantic bridge between them. We hope that by being non-overlapping by design and focused on a single purpose, RDF-STaX will remain useful and relevant in the long term. The three presented use cases, two of which were already realized in practice, are common and well-known to the community, making the ontology all the more relevant.
Another result of this work is the nanopublications dataset, which is a collaborative, living state-of-the-art review of RDF streaming. Nanopublications, in our view, are a natural way to structure the intersubjective scientific discourse. It is possible for example, to comment on other nanopublications, confirm or disagree with them, all in a structured manner. Here we propose to use this powerful tool for a very narrow field of science (RDF streaming), but its possible future applications are certainly much broader Wijkstra et al. (2021). We encourage other researchers to contribute their perspectives to the dataset or to comment on the existing assertions.
As for future work, the theory presented in this contribution notably does not tackle the formal semantics of RDF streams. As with RDF datasets, it is unlikely that there can be single, unifying semantics for streams, but nonetheless, we see this area as worth exploring in the future. The presented survey of RDF streaming is also not a systematic review – it only serves as a basis for the discussion in this work. Producing a systematic review of RDF streaming would be very valuable to the community, and we hope that the proposed taxonomy and the provided nanopublication tooling will make creating it easier.
8 Conclusion
In this work we propose the RDF Stream Taxonomy, which systematizes RDF stream definitions found in the literature. The taxonomy uncovered the variety of RDF streaming in research and practice, highlighting the open challenge of making the different applications compatible with each other. The proposed RDF-STaX ontology embodies the semantic relations between stream types in the taxonomy, and makes the definitions readily applicable. The ontology positions itself as complementary to existing vocabularies, and employs a flexible approach of making subjective statements about the types of RDF streams. The ontology and its documentation are fully open and follow community-developed best practices for FAIR.
The ontology was designed with three use cases in mind, with supporting competency questions for each. Two of these use cases were already implemented – annotating research works in nanopublications, and describing RDF stream types of streaming datasets. The nanopublications dataset is an added contribution of this work, constituting a living state-of-the-art review that anyone can contribute to, using the provided documentation and tooling.
@@ 8 Conclusion
Conceptualization, Piotr Sowiński, Paweł Szmeja and Maria Ganzha; Data curation, Piotr Sowiński; Formal analysis, Piotr Sowiński; Funding acquisition, Maria Ganzha and Marcin Paprzycki; Investigation, Piotr Sowiński; Methodology, Piotr Sowiński and Paweł Szmeja; Project administration, Maria Ganzha and Marcin Paprzycki; Resources, Maria Ganzha and Marcin Paprzycki; Software, Piotr Sowiński; Supervision, Maria Ganzha; Validation, Piotr Sowiński, Paweł Szmeja and Maria Ganzha; Visualization, Piotr Sowiński; Writing – original draft, Piotr Sowiński; Writing – review & editing, Paweł Szmeja, Maria Ganzha and Marcin Paprzycki.
[…]
This research received no external funding.
[…]
Not applicable.
[…]
Not applicable.
[…]
The RDF-STaX ontology introduced in this study (along with the documentation and the nanopublications dataset) is openly available at https://w3id.org/stax/1.1.1 (accessed on 4 June 2024) and in Zenodo at https://zenodo.org/doi/10.5281/zenodo.10072907 (accessed on 4 June 2024), reference number 10072907.
[…]
The authors declare no conflicts of interest.
[…]
Abbreviations The following abbreviations are used in this manuscript:
CI/CD Continuous Integration and Continuous Delivery
CQ Competency Question
DCAT Data Catalog Vocabulary
FAIR Findable, Accessible, Interoperable, Reusable
FOOPS! Ontology Pitfall Scanner for FAIR
GoT Graph of Things
I/O Input/Output
IPSM Inter Platform Semantic Mediator
IRI Internationalized Resource Identifier
LDES Linked Data Event Streams
OOPS! Ontology Pitfall Scanner
OWL Web Ontology Language
RDF Resource Description Framework
RDF-STaX RDF Stream Taxonomy
ROBOT ROBOT is an OBO Tool
RML RDF Mapping Language
RSP RDF Stream Processing
SHACL Shapes Constraint Language
SKOS Simple Knowledge Organization System
SPARQL SPARQL Protocol And RDF Query Language
TA-RDF Time-Annotated RDF
URI Uniform Resource Identifier
W3C World Wide Web Consortium
{adjustwidth}
[…]